OpenStudy (anonymous):

Find the sum. 16 + ( –9 ) = 7 –7 25 –25

OpenStudy (nathan917):

-7 i think.

OpenStudy (anonymous):

can you answer this one to please

OpenStudy (shubhamsrg):

plot 16 in number line, from there, go 9 units back. tell me what you get, its a magic trick! ^_^

OpenStudy (anonymous):

7

OpenStudy (anonymous):

-7 is the answer i know that one i need help one this one Which of the following correctly explains how to model −5 + 3 on a number line? (1 point)Start at zero and move right five units and then left three units to end at –2. Start at zero and move left five units and then right three units to end at −2. Start at zero and move left five units and then right three units to end at 2.

OpenStudy (anonymous):

Which of the following correctly explains how to model −5 + 3 on a number line 1.)Start at zero and move right five units and then left three units to end at –2. 2.)Start at zero and move left five units and then right three units to end at −2. 3.)Start at zero and move left five units and then right three units to end at 2.

OpenStudy (nathan917):

1.

OpenStudy (anonymous):

@engtaha_mg it's -7 i think

OpenStudy (anonymous):

waaaaaaaaaaaaait to answer

OpenStudy (nathan917):

Because -5 would be going right on the number line.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

no no no it is 7

OpenStudy (anonymous):

show me your work @engtaha_mg

OpenStudy (anonymous):

explain 16 + (-9) = 7

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

Start at zero and move right 16 units and then right (-9) units ( it's mean left 9 unit) to end at 7

OpenStudy (nathan917):

My bad it is 7.... *guess that's what i get for answering 4 questions a once...*

OpenStudy (nathan917):

But.. Two like signs become a positive sign Two unlike signs become a negative sign

OpenStudy (nathan917):

so?....

OpenStudy (nathan917):

they are unlike signs

OpenStudy (nathan917):

16−(−9) =16+9 = 7 it is 7.